xx WaveToyFreeF90

WaveToyFreeF90

Abstract

1 Introduction 2 Physical System 3 Numerical Implementation 4 Using This Thorn 4.1 Obtaining This Thorn 4.2 Basic Usage 4.3 Special Behaviour 4.4 Interaction With Other Thorns 4.5 Support and Feedback 5 History 5.1 Thorn Source Code 5.2 Thorn Documentation 5.3 Acknowledgements References

6 Parameters

11 bound 10 Scope: restricted01  KEYWORD

11 Description: Type of boundary condition to use

11 Range 10                01  Default: none
11 none  10 Apply no boundary condition
11 flat  10 Flat (von Neumann, n grad phi = 0) boundary condition
11 static  10 Static (Dirichlet, dphi/dt=0) boundary condition
11 radiation 10 Radiation boundary condition
11 robin  10 Robin (phi(r) = C/r) boundary condition
11 zero  10 Zero (Dirichlet, phi=0) boundary condition

11

7 Interfaces

General Implements: wavetoy

Inherits: grid

Grid Variables

7.0.1 PUBLIC GROUPS

11 x Group Names x 10 x Variable Names x01 Details x 11 x
11  scalarevolve  10                    01  compact 11 0
11                    10  phi  01 description 11 The evolved scalar field
11                    10                    01 dimensions 11 3
11                    10                    01 distribution 11 DEFAULT
11                    10                    01 group type 11 GF
11                    10                    01 timelevels 11 3
11                    10                    01 variable type 11 REAL

11

Uses header: Symmetry.h 8 Schedule This section lists all the variables which are assigned storage by thorn CactusWave/WaveToyFreeF90. Storage can either last for the duration of the run (Always means that if this thorn is activated storage will be assigned, Conditional means that if this thorn is activated storage will be assigned for the duration of the run if some condition is met), or can be turned on for the duration of a schedule function.

Storage x

11 Always:  11 x
11 scalarevolve[3] 11 x

11 x  11 x

11

Scheduled Functions

CCTKxSTARTUP x wavetoyfreef90xstartup x register banner x

11 x 11 Language: 11 fortran
11 x 11 Type:  11 function

11

CCTKxBASEGRID x wavetoyfreef90xinitsymbound x schedule symmetries x

11 x 11 Language: 11 fortran
11 x 11 Type:  11 function

11

CCTKxEVOL x wavetoyfreef90xevolution x evolution of 3d wave equation x

11 x 11 Language: 11 fortran
11 x 11 Sync:  11 scalarevolve
11 x 11 Type:  11 function

CCTKxEVOL x wavetoyfreef90xboundaries x boundaries of 3d wave equation x

11 x 11 After:  11 wavetoyxevolution
11 x 11 Language: 11 fortran
11 x 11 Options: 11 level
11 x 11 Type:  11 function

11

CCTKxEVOL x applybcs x apply boundary conditions x

11 x 11 After: 11 wavetoyxboundaries
11 x 11 Type: 11 group

11

CCTKxPOSTRESTRICT x wavetoyfreef90xboundaries x boundaries of 3d wave equation x

11 x 11 Language: 11 fortran
11 x 11 Options: 11 level
11 x 11 Type:  11 function

11

CCTKxPOSTRESTRICT x applybcs x apply boundary conditions x

11 x 11 After: 11 wavetoyxboundaries
11 x 11 Type: 11 group

11

Aliased Functions x

11 Alias Name: xxxxxxx  11 Function Name:
11 ApplyBCs  11 WaveToyFreeF90xApplyBCs
11 WaveToyFreeF90xBoundaries 11 WaveToyxBoundaries
11 WaveToyFreeF90xEvolution 11 WaveToyxEvolution